OpenStudy (ridwan):

Using the numbers 1, 2, 3, 4 only once and the operations + or × as many times as you like, write an expression that evaluates to 16. Use this expression and any combination of those symbols as many times as you like to write an expression that evaluates to 816.

OpenStudy (ridwan):

I just need the second one, since the first one is 3+1+4 x 2

OpenStudy (gabrian):

hi, i think the key, is the answer of first statement. \[\left[ \left( 3+1+4 \right)2 \right]=16\] call this expresion a; then \[\left( a+1 \right)3a=816\]

OpenStudy (bonnieisflash1.0):

hellur

OpenStudy (joshmorris):

hi